OpenStudy (yusmin):

Can someone please help me? the scale on the map is approximately 5 over 8 inch to 25 miles. If the distance from city A to city B measures 2.5 inches on the​ map, how far apart are the two​ cities?

OpenStudy (mww):

5/8 : 35 miles is your scale of map distance:actual distance As our map distance is 2.5 we need to find the actual distance that would give the same ratio. So we write 5/8 : 35 = 2.5 : x This is equivalent to saying: \[\frac{ (5/8) }{ 35 } = \frac{ 2.5 }{ x }\] Solve for x by cross multiplying and rearranging \[\frac{ 5x }{ 8 } = 2.5 \times 35 = 87.5 \rightarrow x = \frac{ 87.5 \times 8}{ 5 }\] In general you need to find the scaled ratio (given to you) Then any other measurement has to relate to that given scale for which we can write a proportion (i.e. two ratios equalling each other, and solve for the unknown)